Solve for x on the following problems:
x/2 = 7/3
and
3/x = -2/9

Answer by gonzo(654)

problem 1:
x/2 = 7/3
x = (2*7)/3
x = 14/3
problem 2:
3/x = (-2)/9
3 = ((-2)*x)/9
(3*9) = (-2)*x
(3*9)/(-2) = x
27/(-2) = x